Ammonia contains 82.65 % N and 17.65% H2. If the law of constant proportions is true, then the mass of nitrogen required to give 10 g Ammonia will be:
 

Grade:11

1 Answers

Vikas TU
7 years ago
The chemical eqn. would be:
N2  +  H2  ------------>  2NH3
                                 10/17 moles
moles of N2 = 5/17
mass of N2 = 5*28/17 = 8.235 gm Nitrogen